Minutes PSB Upgrade WG Meeting 202, 10th of October 2017

Participants: Fanouria Antoniou, Thomas William Birtwistle, Jean-Baptiste Bonnamy, Jan Borburgh, Julie Coupard, Louis De Mallac, Gian Piero Di Giovanni, Jose Antonio Ferreira Somoza, Alessandro Floriduz, Eve Fortescue-Beck, Robert Froeschl, Georgi Minchev Georgiev, Matthias Haase, Christoph Hessler, Jean-Michel Lacroix, Bettina Mikulec, Antony Newborough, Denis Pazem, Gabor Petrika, Serge Pittet, Theo Prins, Andrea Santamaria Garcia, Frank Schmidt, Jocelyn Tan, Wim Weterings.

Indico event: https://indico.cern.ch/event/669762/


1. Approval of minutes

The minutes of the LIU-PSB WG #201 were approved.

2. Communications

Report on the LIU Project meeting:

3. Follow-up of open actions

M. Buzio (2017-10-02) Provide document describing the specifications of the B-Train system.
  • K. Hanke will contact him.

J-M. Lacroix (2017-09-30) Complete integration of PSB-PS transfer line.

  • Action presented during this meeting and closed.

B. Riffaud (2017-09-30) Provide a proposal for the support of the equipment in the BTP line.

  • Action postponed to 2017-11-31.

G. Petrika (2017-10-10) Prepare a document summarizing the ventilation needs for LIU-PSB (EDMS document 1844566).

  • There is one open question pending, which will be iterated upon next week. The document is to be released in EDMS. Action closed.

G. Petrika (2017-10-10) Verify the consolidation required for building 141.

  • Action closed.

New action: K. Hanke (2017-10-24) Find space to store the materials for the finement cavities installation.

G. Petrika (2017-10-10) release a new version of the EDMS document 1758628, summarizing the cooling needs for LIU-PSB.

  • No new version. Action closed.

G. Petrika (2017-10-10): circulate an EDMS document regarding the new needs of the POPS-B.

  • The document was sent to G. P. Di Giovanni and needs to be circulated for engineering check. Action closed.

L. A. Lopez-Hernandez (2017-09-30) Report about the status of the building 245 modifications.

  • D. Pazem reported that some works are ongoing and will be finished by the end of the year. Action closed.

Assigned to Due date Description State Result  
J-M.Lacroix 2017-09-30 Complete integration of PSB-PS transfer line edit
G.Petrika 2017-10-10 Prepare a document summarizing the ventilation needs for LIU-PSB. edit
G.Petrika 2017-10-10 Verify the consolidation required for building 141. edit
G.P.Di.Giovanni 2018-02-06 Find space to store the materials for the finement cavities installation. M. Haase confirmed that the space in B192 will be free at the end of March. edit
G.Petrika 2017-10-10 release a new version of the EDMS document 1758628, summarizing the cooling needs for LIU-PSB. edit
G.Petrika 2017-10-10 Circulate an EDMS document regarding the new needs of the POPS-B edit
L.A.Lopez-Hernandez 2017-09-30 Report about the status of the building 245 modifications.   edit

4. BTP Line Status (Jean-Michel Lacroix)

  • There were some discussions regarding the beam stopper since the team did not have the same values as T. Dobers. There is still a difference concerning the wide band pick-up. A 3D scan was made and the data is available. It is assumed that the scan data are correct.
  • The support of the dipoles is still to be reviewed, which will be done by B. Riffaud.
  • There were no problems with the new BPMs and the BT line integration.

New Action:
Assigned to Due date Description State Result  
T.Dobers 2017-11-21 Change the value of the wide band pick-up in Geode to the one given by the 3D scan. edit

  • It needs to be confirmed that B. Riffaud has all the information to continue working on the dipole supports. He should present in the next ICL meeting to define the next actions to be taken. J. Coupard will talk to J. P. Corso and make sure the topic is on the agenda for the next meeting.
  • A wall needs to be carved a few centimeters to make the magnets fit.

5. AOB

Cabling for YETS 17-18 in PSB (Georgi Minchev Georgiev)
  • The execution of the cabling operations is estimated to be 6.5 weeks.
  • Concerning the budget, the engagement will be in 2017 and paid in 2018.
  • The cabling for the Q-strip will not be done this year but in LS2.
  • The POPS-B cabling will be done by a different team running in parallel, and the budget is associated with the TE-EPC budget code.

New BPM in the BI line (Christoph Hessler)
  • They had a first look at integration and they will place the new BPM as close as possible to the SMV10 tank, upstream.
  • They will place the BPM upstream of the BTV30, cutting a portion of the vacuum chamber.
  • All the changes will be discussed in the ICL meeting next week.
  • A space reservation request for the BPM will be placed.

  • J. A. Ferreira Somoza commented that it might be easier to produce another vacuum chamber instead of cutting the current one. J-M. Lacroix said that, indeed, they will need to cut the chamber in other additional points if that option is chosen.
  • J. Coupard proposed to already prepare the ECR concerning the BPM. C. Hessler answered that it will take some more time since the technical details are still not ready.

ECR for the coming YETS and LS2 (Julie Coupard)
  • The deadlines to submit the ECRs for the YETS has already passed.
  • The deadline to provide the ECRs for the LS2 document is March 2018, although it is recommended to send them before the end of 2017.
  • Concerning the list of ECRs to be provided by the LIU-PSB WG, it is important to provide as soon as possible the one describing the SEMGRID installation since they will be installed during the YETS.
  • The ECRs need to be validated by integration before being sent for approval.

6. Tour de table

  • J. A. Ferreira Somoza reported that they will provide a proposal for the modification of the vacuum chamber for the BPM in the BI line.
  • J. Coupard commented that they are currently working on the LS2 schedule.
  • L. De Mallac said that discussions are ongoing concerning the injection Q-strips. There is a small working group on this topic.
  • G. M. Georgiev commented that the measures of the pits are very important to continue the integration and determine if the blocks should be removed or not. It will be followed-up by R. Froeschl.
  • T. Birtwistle reported that T. Prins will be working on the post-LS2 layout database.
  • M. Haase reported that they assembled all the finemet cavities. The first test with a block of 6 cells gave good results. There are some problems with the vacuum chambers.
  • F. Schmidt commented that there were some fruitful discussions during the space-charge workshop and that the computational tools (PyOrbit) are ready for the PSB.

Next LIU-PSB meeting: 24th of October on the TFB status by Alfred Blas.

-- AndreaSantamariaGarcia - 2017-10-11

Edit | Attach | Watch | Print version | History: r12 < r11 < r10 < r9 < r8 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r12 - 2018-02-19 - unknown
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSBUpgrade All webs login

This site is powered by the TWiki collaboration platform Powered by PerlWeb page setup by Thomas Hermanns